# Bondal-Orlov’s reconstruction theorem in noncommutative projective geometry

### v1

@@ 1. Introduction
If $-\otimes\mathcal{K}_{A},-\otimes\mathcal{K}_{B}$ are AZ-(anti-)ample (see 4.16), then
@@
Corollary 1.6 (= 4.21).
Let $A,B$ be noetherian AS-regular algebras over $A_{0},B_{0}$, respectively (see 4.19). Then,
@@ 2.2. Noncommutative projective geometry
    the natural morphisms $A\rightarrow\mathbf{R}\operatorname{Hom}_{A}(R,R)$ and $A^{\operatorname{op}}\rightarrow\mathbf{R}\operatorname{Hom}_{A^{\operatorname{op}}}(R,R)$ are isomorphisms in $D(\operatorname{Gr}(A^{\operatorname{en}}))$.
$R$ is called balanced if $\mathbf{R}\Gamma_{\mathfrak{m}_{A}}(R)\simeq A^{\prime}$ and $\mathbf{R}\Gamma_{\mathfrak{m}_{A^{\operatorname{op}}}}(R)\simeq A^{\operatorname{op}^{\prime}}$ in $D(\operatorname{Gr}(A^{\operatorname{en}}))$.
@@
$\pi_{A}(M)$ $\mapsto\pi(M\otimes_{A}^{\mathbf{L}}R_{A})[-1]$
@@ 4. Bondal-Orlov’s reconstruction theorem for noncommutative projective schemes
For simplicity, we set $n=\operatorname{gl.dim}(\operatorname{qgr}(A))$ and $m=\operatorname{gl.dim}(\operatorname{qgr}(B))$. Let $F:D^{b}(\operatorname{qgr}(A))\rightarrow D^{b}(\operatorname{qgr}(B))$ be an equivalence. Then, from (3) of 3.10, we have $F\simeq\Phi_{\mathcal{F}}$ for some $\mathcal{F}\in D(\operatorname{QGr}(A^{\operatorname{op}}\otimes_{k}B))$ and $\mathcal{F}$ is unique up to quasi-isomorphism. Let $G:D^{b}(\operatorname{qgr}(B))\rightarrow D^{b}(\operatorname{qgr}(A))$ be the quasi-inverse of $F$. Then, $G\simeq\Phi_{\mathcal{G}}$ for some $\mathcal{G}\in D(\operatorname{QGr}(B^{\operatorname{op}}\otimes_{k}A))$ and $\mathcal{G}$ is unique up to quasi-isomorphism. Note that when $n=0$ or $m=0$, (3) of 3.10 cannot be applied to at least one of $F$ and $G$. However, because we have an equivalence between $D^{b}_{\operatorname{dg}}(\operatorname{qgr}(A)$) and $D^{b}_{\operatorname{dg}}(\operatorname{qgr}(B))$ from the uniqueness of dg enhancements and obtain an equivalence between $D_{\operatorname{dg}}(\operatorname{QGr}(A))$ and $D_{\operatorname{dg}}(\operatorname{QGr}(B))$ from ($\dagger$3), so we can also take Fourier-Mukai functors as an equivalence $F:D^{b}(\operatorname{qgr}(A))\rightarrow D^{b}(\operatorname{qgr}(B))$ and a quasi-inverse $G$ of $F$ (cf. [ballard2021kernels, Corollary 4.18]). Then, $G$ is a left adjoint of $F$. Moreover, $H:=\operatorname{\mathcal{S}}_{D^{b}(\operatorname{qgr}(A))}\circ G\circ\operatorname{\mathcal{S}}_{D^{b}(\operatorname{qgr}(B))}^{-1}$ is a right adjoint of $F$ ([huybrechts2006fourier, Remark 3.31]), where $\operatorname{\mathcal{S}}_{D^{b}(\operatorname{qgr}(A))}$ (resp. $\operatorname{\mathcal{S}}_{D^{b}(\operatorname{qgr}(B))}$) is the Serre functor of $D^{b}(\operatorname{qgr}(A))$ (resp. $D^{b}(\operatorname{qgr}(B))$). Thus, we have $G\simeq H$. In addtition, $\operatorname{\mathcal{S}}_{D^{b}(\operatorname{qgr}(A))}$ and $\operatorname{\mathcal{S}}_{D^{b}(\operatorname{qgr}(B))}$ are given by $-\otimes\pi_{A}\left(H^{-(n+1)}(R_{A})\right)[n]$ and $-\otimes\pi_{B}\left(H^{-(m+1)}(R_{B})\right)[m]$, respectively by our assumption. So, we have
@@
′′$\mathcal{E}^{\prime\prime}=\pi_{A\otimes_{k}B^{\operatorname{op}}}(\mathbf{R}Q_{B^{\operatorname{op}}\otimes_{k}B}(H^{-(m+1)}(R_{B}^{\operatorname{bi}})^{\vee})\otimes_{\mathcal{B}}^{\mathbf{L}}\mathbf{R}\omega_{A\otimes_{k}B^{\operatorname{op}}}(\mathcal{E}^{\prime}))[-m],$
where $(-)^{\vee}=\operatorname{Hom}_{B}(-,B)$. By using 4.7 again, the kernel $\mathcal{E}^{\prime\prime\prime}$ of $H$ is given by
′′′$\mathcal{E}^{\prime\prime\prime}=\pi_{A\otimes_{k}B^{\operatorname{op}}}(\mathbf{R}\omega_{A\otimes_{k}B^{\operatorname{op}}}(\mathcal{E}^{\prime\prime})\otimes_{\mathcal{A}}^{\mathbf{L}}\mathbf{R}Q_{A^{\operatorname{op}}\otimes_{k}A}(H^{-(n+1)}(R_{A}^{\operatorname{bi}})))[n].$
@@
$=\pi_{A\otimes_{k}B^{\operatorname{op}}}(\mathbf{R}Q_{B^{\operatorname{op}}\otimes_{k}B}(M_{B}^{\operatorname{bi}{\vee}})\otimes_{\mathcal{B}}^{\mathbf{L}}\mathbf{R}\omega_{B^{\operatorname{op}}\otimes_{k}A}(\mathcal{E}^{\prime})\otimes_{\mathcal{A}}^{\mathbf{L}}\mathbf{R}Q_{A^{\operatorname{op}}\otimes_{k}A}(M_{A}^{\operatorname{bi}}))[n-m]$
$=\pi_{A\otimes_{k}B^{\operatorname{op}}}(\mathbf{R}^{\prime}Q_{B^{\operatorname{op}}}(M_{B}^{\operatorname{bi}\vee})\otimes_{\mathcal{B}}^{\mathbf{L}}\mathbf{R}\omega_{B^{\operatorname{op}}\otimes_{k}A}(\mathcal{E}^{\prime})\otimes_{\mathcal{A}}^{\mathbf{L}}\mathbf{R}^{\prime}Q_{A}(M_{A}^{\operatorname{bi}}))[n-m]$
$=\pi_{A\otimes_{k}B^{\operatorname{op}}}(\mathbf{R}^{\prime}Q_{B^{\operatorname{op}}}(M_{B}^{\operatorname{bi}\vee}\otimes_{\mathcal{B}}^{\mathbf{L}}\mathbf{R}\omega_{B^{\operatorname{op}}\otimes_{k}A}(\mathcal{E}^{\prime})\otimes_{\mathcal{A}}^{\mathbf{L}}\mathbf{R}^{\prime}Q_{A}(M_{A}^{\operatorname{bi}})))[n-m]$
$=\pi_{A\otimes_{k}B^{\operatorname{op}}}((\mathbf{R}^{\prime}Q_{B^{\operatorname{op}}}\circ\mathbf{R}^{\prime}Q_{A})(M_{B}^{\operatorname{bi}\vee}\otimes_{\mathcal{B}}^{\mathbf{L}}\mathbf{R}\omega_{B^{\operatorname{op}}\otimes_{k}A}(\mathcal{E}^{\prime})\otimes_{\mathcal{A}}^{\mathbf{L}}\mathbf{R}^{\prime}Q_{A}(M_{A}^{\operatorname{bi}})))[n-m]$
$=(\pi_{A\otimes_{k}B^{\operatorname{op}}}\circ\mathbf{R}Q_{A\otimes_{k}B^{\operatorname{op}}})(M_{B}^{\operatorname{bi}\vee}\otimes_{\mathcal{B}}^{\mathbf{L}}\mathbf{R}\omega_{B^{\operatorname{op}}\otimes_{k}A}(\mathcal{E}^{\prime})\otimes_{\mathcal{A}}^{\mathbf{L}}M_{A}^{\operatorname{bi}})[n-m]$
$=\pi_{A\otimes_{k}B^{\operatorname{op}}}(M_{B}^{\operatorname{bi}\vee}\otimes_{\mathcal{B}}^{\mathbf{L}}\mathbf{R}\omega_{B^{\operatorname{op}}\otimes_{k}A}(\mathcal{E}^{\prime})\otimes_{\mathcal{A}}^{\mathbf{L}}M_{A}^{\operatorname{bi}})[n-m]$
$=\pi_{A\otimes_{k}B^{\operatorname{op}}}(M_{B}^{\operatorname{bi}\vee}\otimes_{\mathcal{B}}\mathbf{R}\omega_{B^{\operatorname{op}}\otimes_{k}A}(\mathcal{E}^{\prime})\otimes_{\mathcal{A}}M_{A}^{\operatorname{bi}})[n-m].$
@@
This functor is an equivalence since $\Phi_{\mathcal{F}}$ and $\Phi^{\prime}_{\mathcal{G}}$ are equivalences (in detail, see 4.15).
As in the proof of 4.10, we put $M_{A}=H^{-(n+1)}(R_{A})$ and $M_{B}=H^{-(m+1)}(R_{B})$, where $n,m$ are the global dimensions of $\operatorname{qgr}(A),\operatorname{qgr}(B)$. We also put $\mathcal{M}_{l}=\tilde{\Phi}_{\mathcal{H}}(\pi_{A\otimes_{k}A^{\operatorname{op}}}(M_{A}^{\otimes l\operatorname{bi}}))$. Then, we show that
@@
We see that the kernel of $\Phi_{\mathcal{F}}\circ\Phi_{\pi_{A\otimes_{k}A^{\operatorname{op}}}\left(M_{A}^{\otimes l\operatorname{bi}}\right)}\circ\Phi_{\mathcal{G}}$ is isomorphic to $\mathcal{M}_{l}$. Actually, from 4.4, 4.5, 4.7 and 4.14, we have
@@
Note that $\Phi_{\pi_{A\otimes_{k}A^{\operatorname{op}}}\left(M_{A}^{\otimes l\operatorname{bi}}\right)}\simeq\mathcal{S}_{A}^{l}[-ln]$, where $\mathcal{S}_{A}$ is the Serre functor of $D^{b}(\operatorname{qgr}(A))$. So, since Serre functors commute with equivalences, we have $\Phi_{\mathcal{M}_{l}}\simeq\mathcal{S}_{B}[-ln]$. Moreover, by 4.10 and the uniqueness of Fourier-Mukai kernels ((3) of 3.10), we have $\mathcal{M}_{l}\simeq\pi_{B\otimes_{k}B^{\operatorname{op}}}(M_{B}^{\otimes l\operatorname{bi}})$. Thus, for all $l\in\mathbb{Z}$, we have $\tilde{\Phi}_{\mathcal{H}}(\pi_{A\otimes_{k}A^{\operatorname{op}}}(M_{A}^{\otimes l\operatorname{bi}}))\simeq\pi_{B\otimes_{k}B^{\operatorname{op}}}(M_{B}^{\otimes l\operatorname{bi}})$.
@@
$\widetilde{\Phi}_{\mathcal{H}^{\prime}}:D(\operatorname{QBiGr}(B\otimes_{k}B^{\operatorname{op}}))\rightarrow D(\operatorname{QBiGr}(A\otimes_{k}A^{\operatorname{op}}))$
@@
    for every epimorphism $f:\mathcal{M}\rightarrow\mathcal{N}$ in $\mathcal{C}$, there exists $n_{0}$ such that the natural map $\operatorname{Hom}_{\mathcal{C}}(\mathcal{O},s^{n}(f)):\operatorname{Hom}_{\mathcal{C}}(\mathcal{O},s^{n}(\mathcal{M}))\rightarrow\operatorname{Hom}_{\mathcal{C}}(\mathcal{O},s^{n}(\mathcal{N}))$ for all $n\geq n_{0}$ is surjective.
@@
We assume that $-\otimes\mathcal{K}_{A},-\otimes\mathcal{K}_{B}$ are AZ-anti-ample. Then, from 4.12, we have
@@
by 4.17. In the case that $-\otimes\mathcal{K}_{A},-\otimes\mathcal{K}_{B}$ are AZ-ample, we can prove the theorem in a similar way.
In the following, we give an application of the argument of 4.18 for AS-regular algebras. We recall the definitions of AS-regular algebras and AS-Gorenstein algebras below.
@@
The application of 4.18 is the following corollary.

### v2

@@ 1. Introduction
If $-\otimes\mathcal{K}_{A},-\otimes\mathcal{K}_{B}$ are AZ-(anti-)ample (see 4.18), then
@@
Corollary 1.6 (= 4.23).
Let $A,B$ be noetherian AS-regular algebras over $A_{0},B_{0}$, respectively (see 4.21). Then,
@@ 2.2. Noncommutative projective geometry
    the natural morphisms $A\rightarrow\mathbf{R}\operatorname{Hom}_{A}(R,R)$ and $A\rightarrow\mathbf{R}\operatorname{Hom}_{A^{\operatorname{op}}}(R,R)$ are isomorphisms in $D(\operatorname{Gr}(A^{\operatorname{en}}))$.
$R$ is called balanced if $\mathbf{R}\Gamma_{\mathfrak{m}_{A}}(R)\simeq A^{\prime}$ and $\mathbf{R}\Gamma_{\mathfrak{m}_{A^{\operatorname{op}}}}(R)\simeq A^{\prime}$ in $D(\operatorname{Gr}(A^{\operatorname{en}}))$.
@@
$\pi_{A}(M)$ $\mapsto\pi_{A}(M\otimes_{A}^{\mathbf{L}}R_{A})[-1]$
@@ 4. Bondal-Orlov’s reconstruction theorem for noncommutative projective schemes
Lemma 4.10.
Let $A$ be a locally finite noetherian $\mathbb{N}$-graded $k$-algebra with a balanced dualizing complex. Let $M\in D(\operatorname{Gr}(A^{\operatorname{en}}))$. Then, the functor
$\pi_{A}(-\otimes^{\mathbf{L}}_{A}M):D(\operatorname{QGr}A)\rightarrow D(\operatorname{QGr}A),\quad\pi_{A}(N)\mapsto\pi_{A}(N\otimes^{\mathbf{L}}_{A}M)$
is naturally isomorphic to $\Phi_{\pi_{A^{\operatorname{en}}}(M^{\operatorname{bi}})}$.
Proof.
We show that
$\pi_{A}(\mathbf{R}Q_{A}(N)\otimes^{\mathbf{L}}_{\mathcal{A}}\mathbf{R}Q_{A^{en}}(M^{\operatorname{bi}}))\simeq\pi_{A}(N\otimes^{\mathbf{L}}_{A}M).$
Firstly, because
$\mathbf{R}\tau_{A}(N)\otimes^{\mathbf{L}}_{\mathcal{A}}\mathbf{R}Q_{A^{en}}(M^{\operatorname{bi}})\simeq\mathbf{R}\tau_{A}(N)\otimes_{\mathcal{A}}^{\mathbf{L}}\mathbf{R}^{\prime}Q_{A^{\operatorname{op}}}\mathbf{R}^{\prime}Q_{A}(M^{\operatorname{bi}})$
is acyclic from the locally finite version of [ballard2021kernels, Proposition 4.6], we have
$\mathbf{R}Q_{A}(N)\otimes_{\mathcal{A}}^{\mathbf{L}}\mathbf{R}Q_{A^{en}}(M^{\operatorname{bi}})$ $\simeq N\otimes_{\mathcal{A}}^{\mathbf{L}}\mathbf{R}Q_{A^{en}}(M^{\operatorname{bi}})$
$\simeq N\otimes_{\mathcal{A}}^{\mathbf{L}}\mathbf{R}^{\prime}Q_{A}(M^{\operatorname{bi}})$
$\simeq\mathbf{R}Q_{A}(N\otimes_{\mathcal{A}}^{\mathbf{L}}M^{\operatorname{bi}}).$
Note that the second isomorphism is obtained by using 4.6 and the third isomorphism is obtained by using 4.5. So, we show that
$\pi_{A}(N\otimes_{\mathcal{A}}^{\mathbf{L}}M^{\operatorname{bi}})\simeq\pi_{A}(N\otimes_{A}^{\mathbf{L}}M).$
Moreover, since $\operatorname{Gr}(A)$ is equivalent to the category of right modules of the $\mathbb{Z}$-algebra $\mathcal{A}$, it is enough to show that $N\otimes_{\mathcal{A}}M^{\operatorname{bi}}$ and $N\otimes_{A}M$ are isomorphic as modules of the $\mathbb{Z}$-algebra $\mathcal{A}$ for any $N\in\operatorname{Gr}(A)$ and $M\in\operatorname{Gr}(A^{\operatorname{en}})$. We can easily check this from the definition of the tensor product of modules of a $\mathbb{Z}$-algebra. ∎
@@
For simplicity, we set $n=\operatorname{gl.dim}(\operatorname{qgr}(A))$ and $m=\operatorname{gl.dim}(\operatorname{qgr}(B))$. Let $F:D^{b}(\operatorname{qgr}(A))\rightarrow D^{b}(\operatorname{qgr}(B))$ be an equivalence. Then, from (3) of 3.10, we have $F\simeq\Phi_{\mathcal{F}}$ for some $\mathcal{F}\in D(\operatorname{QGr}(A^{\operatorname{op}}\otimes_{k}B))$ and $\mathcal{F}$ is unique up to quasi-isomorphism. Let $G:D^{b}(\operatorname{qgr}(B))\rightarrow D^{b}(\operatorname{qgr}(A))$ be the quasi-inverse of $F$. Then, $G\simeq\Phi_{\mathcal{G}}$ for some $\mathcal{G}\in D(\operatorname{QGr}(B^{\operatorname{op}}\otimes_{k}A))$ and $\mathcal{G}$ is unique up to quasi-isomorphism. Note that when $n=0$ or $m=0$, (3) of 3.10 cannot be applied to at least one of $F$ and $G$. However, because we have an equivalence between $D^{b}_{\operatorname{dg}}(\operatorname{qgr}(A)$) and $D^{b}_{\operatorname{dg}}(\operatorname{qgr}(B))$ from the uniqueness of dg enhancements and obtain an equivalence between $D_{\operatorname{dg}}(\operatorname{QGr}(A))$ and $D_{\operatorname{dg}}(\operatorname{QGr}(B))$ from ($\dagger$3), so we can also take Fourier-Mukai functors as an equivalence $F:D^{b}(\operatorname{qgr}(A))\rightarrow D^{b}(\operatorname{qgr}(B))$ and a quasi-inverse $G$ of $F$ (cf. [ballard2021kernels, Corollary 4.18]). Then, $G$ is a left adjoint of $F$. Moreover, $H:=\operatorname{\mathcal{S}}_{D^{b}(\operatorname{qgr}(A))}\circ G\circ\operatorname{\mathcal{S}}_{D^{b}(\operatorname{qgr}(B))}^{-1}$ is a right adjoint of $F$ ([huybrechts2006fourier, Remark 3.31]), where $\operatorname{\mathcal{S}}_{D^{b}(\operatorname{qgr}(A))}$ (resp. $\operatorname{\mathcal{S}}_{D^{b}(\operatorname{qgr}(B))}$) is the Serre functor of $D^{b}(\operatorname{qgr}(A))$ (resp. $D^{b}(\operatorname{qgr}(B))$). Thus, we have $G\simeq H$. In addtition, $\operatorname{\mathcal{S}}_{D^{b}(\operatorname{qgr}(A))}$ and $\operatorname{\mathcal{S}}_{D^{b}(\operatorname{qgr}(B))}$ are given by $-\otimes_{A}\pi_{A}\left(H^{-(n+1)}(R_{A})\right)[n]$ and $-\otimes_{B}\pi_{B}\left(H^{-(m+1)}(R_{B})\right)[m]$, respectively by our assumption (in detail, see 4.12). So, from 4.10, we have
@@
′′$\mathcal{E}^{\prime\prime}\simeq\pi_{A\otimes_{k}B^{\operatorname{op}}}(\mathbf{R}Q_{B^{\operatorname{op}}\otimes_{k}B}(H^{-(m+1)}(R_{B}^{\operatorname{bi}})^{\vee})\otimes_{\mathcal{B}}^{\mathbf{L}}\mathbf{R}\omega_{A\otimes_{k}B^{\operatorname{op}}}(\mathcal{E}^{\prime}))[-m],$
where $(-)^{\vee}=\operatorname{Hom}_{B}(-,B)$ (in detail, see 4.12). By using 4.7 again, the kernel $\mathcal{E}^{\prime\prime\prime}$ of $H$ is given by
′′′$\mathcal{E}^{\prime\prime\prime}\simeq\pi_{A\otimes_{k}B^{\operatorname{op}}}(\mathbf{R}\omega_{A\otimes_{k}B^{\operatorname{op}}}(\mathcal{E}^{\prime\prime})\otimes_{\mathcal{A}}^{\mathbf{L}}\mathbf{R}Q_{A^{\operatorname{op}}\otimes_{k}A}(H^{-(n+1)}(R_{A}^{\operatorname{bi}})))[n].$
@@
$\simeq\pi_{A\otimes_{k}B^{\operatorname{op}}}(\mathbf{R}Q_{B^{\operatorname{op}}\otimes_{k}B}(M_{B}^{\operatorname{bi}{\vee}})\otimes_{\mathcal{B}}^{\mathbf{L}}\mathbf{R}\omega_{B^{\operatorname{op}}\otimes_{k}A}(\mathcal{E}^{\prime})\otimes_{\mathcal{A}}^{\mathbf{L}}\mathbf{R}Q_{A^{\operatorname{op}}\otimes_{k}A}(M_{A}^{\operatorname{bi}}))[n-m]$
$\simeq\pi_{A\otimes_{k}B^{\operatorname{op}}}(\mathbf{R}^{\prime}Q_{B^{\operatorname{op}}}(M_{B}^{\operatorname{bi}\vee})\otimes_{\mathcal{B}}^{\mathbf{L}}\mathbf{R}\omega_{B^{\operatorname{op}}\otimes_{k}A}(\mathcal{E}^{\prime})\otimes_{\mathcal{A}}^{\mathbf{L}}\mathbf{R}^{\prime}Q_{A}(M_{A}^{\operatorname{bi}}))[n-m]$
$\simeq\pi_{A\otimes_{k}B^{\operatorname{op}}}(\mathbf{R}^{\prime}Q_{B^{\operatorname{op}}}(M_{B}^{\operatorname{bi}\vee}\otimes_{\mathcal{B}}^{\mathbf{L}}\mathbf{R}\omega_{B^{\operatorname{op}}\otimes_{k}A}(\mathcal{E}^{\prime})\otimes_{\mathcal{A}}^{\mathbf{L}}\mathbf{R}^{\prime}Q_{A}(M_{A}^{\operatorname{bi}})))[n-m]$
$\simeq\pi_{A\otimes_{k}B^{\operatorname{op}}}((\mathbf{R}^{\prime}Q_{B^{\operatorname{op}}}\circ\mathbf{R}^{\prime}Q_{A})(M_{B}^{\operatorname{bi}\vee}\otimes_{\mathcal{B}}^{\mathbf{L}}\mathbf{R}\omega_{B^{\operatorname{op}}\otimes_{k}A}(\mathcal{E}^{\prime})\otimes_{\mathcal{A}}^{\mathbf{L}}\mathbf{R}^{\prime}Q_{A}(M_{A}^{\operatorname{bi}})))[n-m]$
$\simeq(\pi_{A\otimes_{k}B^{\operatorname{op}}}\circ\mathbf{R}Q_{A\otimes_{k}B^{\operatorname{op}}})(M_{B}^{\operatorname{bi}\vee}\otimes_{\mathcal{B}}^{\mathbf{L}}\mathbf{R}\omega_{B^{\operatorname{op}}\otimes_{k}A}(\mathcal{E}^{\prime})\otimes_{\mathcal{A}}^{\mathbf{L}}M_{A}^{\operatorname{bi}})[n-m]$
$\simeq\pi_{A\otimes_{k}B^{\operatorname{op}}}(M_{B}^{\operatorname{bi}\vee}\otimes_{\mathcal{B}}^{\mathbf{L}}\mathbf{R}\omega_{B^{\operatorname{op}}\otimes_{k}A}(\mathcal{E}^{\prime})\otimes_{\mathcal{A}}^{\mathbf{L}}M_{A}^{\operatorname{bi}})[n-m]$
$\simeq\pi_{A\otimes_{k}B^{\operatorname{op}}}(M_{B}^{\operatorname{bi}\vee}\otimes_{\mathcal{B}}\mathbf{R}\omega_{B^{\operatorname{op}}\otimes_{k}A}(\mathcal{E}^{\prime})\otimes_{\mathcal{A}}M_{A}^{\operatorname{bi}})[n-m].$
@@
Remark 4.12.
In this remark, we give a detailed description about the Serre functor of $D^{b}(\operatorname{qgr}(A))$ when $A$ is a locally finite noetherian $\mathbb{N}$-graded $k$-algebra with a balanced dualizing complex $R_{A}$ and $\operatorname{qgr}(A)$ has a canonical bimodule $\mathcal{K}_{A}$.
Firstly, we recall that the Serre functor $\mathcal{S}_{D^{b}(\operatorname{qgr}(A))}$ of $D^{b}(\operatorname{qgr}(A))$ is given by
$\mathcal{S}_{D^{b}(\operatorname{qgr}(A))}(\pi_{A}(M))\simeq\pi_{A}(M\otimes^{\mathbf{L}}_{A}R_{A})[-1]\simeq\pi_{A}(M)\otimes\mathcal{K}_{A}[n],$
where $n$ is the global dimension of $\operatorname{qgr}(A)$. In particular, because we have $\pi_{A}(N\otimes^{\mathbf{L}}_{A}R_{A})\simeq\pi_{A}(N)\otimes\mathcal{K}_{A}[n+1]\in\operatorname{qgr}(A)[n+1]$ for all $N\in\operatorname{gr}(A)$, we have $\pi_{A}(R_{A})\simeq\pi_{A}(M_{A})[n+1]\in D^{b}(\operatorname{qgr}(A))$ and
$\mathcal{S}_{D^{b}(\operatorname{qgr}(A))}(\pi_{A}(N))\simeq\pi_{A}(N\otimes^{\mathbf{L}}_{A}M_{A})[n]\simeq\pi_{A}(N\otimes_{A}M_{A})[n],$
where we put $M_{A}=H^{-(n+1)}(R_{A})$ for simplicity. Note that from these isomorphisms, we obtain the well-definedness of $\pi_{A}(-\otimes_{A}M_{A})[n]$ as a endofunctor of $D^{b}(\operatorname{qgr}(A))$.
On the other hand, the inverse of the Serre functor $\mathcal{S}_{D^{b}(\operatorname{qgr}(A))}$ is given by
$\mathcal{S}_{D^{b}(\operatorname{qgr}(A))}^{-1}(\pi_{A}(N))\simeq\pi_{A}(\mathbf{R}\operatorname{Hom}_{A}(R_{A},\pi_{A}(N)))[1]$
from [de2004ideal, Proposition A.3]. Then, we have a natural morphism
$\pi_{A}(N\otimes_{A}^{\mathbf{L}}\mathbf{R}\operatorname{Hom}_{A}(R_{A},A))\rightarrow\pi_{A}(\mathbf{R}\operatorname{Hom}_{A}(R_{A},N)).$
This morphism is an isomorphism because $D^{b}(\operatorname{qgr}(A))$ is classically generated by $\{\mathcal{O}_{A}(m)\}_{m\in\mathbb{Z}}$ ([bondal2002generators, Lemma 4.2.2]) and we can apply [yekutieli2019derived, Propostion 5.3.22]. So, we have
$\mathcal{S}_{D^{b}(\operatorname{qgr}(A))}^{-1}(\pi_{A}(N))$ $\simeq\pi_{A}(N\otimes_{A}^{\mathbf{L}}\mathbf{R}\operatorname{Hom}_{A}(R_{A},A))[1]$
$\simeq\pi_{A}(N\otimes_{A}^{\mathbf{L}}\mathbf{R}\operatorname{Hom}_{A}(M_{A},A))[1]$
$\simeq\pi_{A}(N\otimes^{\mathbf{L}}_{A}M_{A}^{\vee})[1]$
$\simeq\pi_{A}(N\otimes_{A}M_{A}^{\vee})[1],$
where $M_{A}^{\vee}=\operatorname{Hom}_{A}(M_{A},A)$. In addition, we have the following isomorphisms in $D(\operatorname{QGr}(A^{\operatorname{en}}))$:
$\pi_{A^{\operatorname{en}}}(R_{A}\otimes_{A}^{\mathbf{L}}\mathbf{R}\operatorname{Hom}_{A}(R_{A},A))$ $\simeq\pi_{A^{\operatorname{en}}}(M_{A}\otimes_{A}M_{A}^{\vee}),$
$\pi_{A^{\operatorname{en}}}(\mathbf{R}\operatorname{Hom}_{A}(R_{A},R_{A})\otimes_{A}^{\mathbf{L}}R_{A})$ $\simeq\pi_{A^{\operatorname{en}}}(M_{A}^{\vee}\otimes_{A}M_{A}),$
which are used in the last isomorphism of the arrangement of $\mathcal{E}^{\prime\prime\prime}$ in the proof of 4.11.
@@
This functor is an equivalence since $\Phi_{\mathcal{F}}$ and $\Phi^{\prime}_{\mathcal{G}}$ are equivalences (in detail, see 4.17).
As in the proof of 4.11, we put $M_{A}=H^{-(n+1)}(R_{A})$ and $M_{B}=H^{-(m+1)}(R_{B})$, where $n,m$ are the global dimensions of $\operatorname{qgr}(A),\operatorname{qgr}(B)$. We also put $\mathcal{M}_{l}=\tilde{\Phi}_{\mathcal{H}}(\pi_{A\otimes_{k}A^{\operatorname{op}}}(M_{A}^{\otimes l\operatorname{bi}}))$. Then, we show that
@@
We see that the kernel of $\Phi_{\mathcal{F}}\circ\Phi_{\pi_{A\otimes_{k}A^{\operatorname{op}}}\left(M_{A}^{\otimes l\operatorname{bi}}\right)}\circ\Phi_{\mathcal{G}}$ is isomorphic to $\mathcal{M}_{l}$. Actually, from 4.4, 4.5, 4.7 and 4.16, we have
@@
Note that $\Phi_{\pi_{A\otimes_{k}A^{\operatorname{op}}}\left(M_{A}^{\otimes l\operatorname{bi}}\right)}\simeq\mathcal{S}_{A}^{l}[-ln]$, where $\mathcal{S}_{A}$ is the Serre functor of $D^{b}(\operatorname{qgr}(A))$. So, since Serre functors commute with equivalences, we have $\Phi_{\mathcal{M}_{l}}\simeq\mathcal{S}_{B}[-ln]$. Moreover, by 4.11 and the uniqueness of Fourier-Mukai kernels ((3) of 3.10), we have $\mathcal{M}_{l}\simeq\pi_{B\otimes_{k}B^{\operatorname{op}}}(M_{B}^{\otimes l\operatorname{bi}})$. Thus, for all $l\in\mathbb{Z}$, we have $\tilde{\Phi}_{\mathcal{H}}(\pi_{A\otimes_{k}A^{\operatorname{op}}}(M_{A}^{\otimes l\operatorname{bi}}))\simeq\pi_{B\otimes_{k}B^{\operatorname{op}}}(M_{B}^{\otimes l\operatorname{bi}})$.
@@
$\widetilde{\Phi}_{\mathcal{H}^{\prime}}:D(\operatorname{QBiGr}(B\otimes_{k}B^{\operatorname{op}}))\rightarrow D(\operatorname{QBiGr}(A\otimes_{k}A^{\operatorname{op}})),$
@@
    for every epimorphism $f:\mathcal{M}\rightarrow\mathcal{N}$ in $\mathcal{C}$, there exists $n_{0}$ such that the natural map $\operatorname{Hom}_{\mathcal{C}}(s^{-n}(\mathcal{O}),\mathcal{M})\rightarrow\operatorname{Hom}_{\mathcal{C}}(s^{-n}(\mathcal{O}),\mathcal{N})$ for all $n\geq n_{0}$ is surjective.
@@
We assume that $-\otimes\mathcal{K}_{A},-\otimes\mathcal{K}_{B}$ are AZ-anti-ample. Then, from 4.14, we have
@@
by 4.19. In the case that $-\otimes\mathcal{K}_{A},-\otimes\mathcal{K}_{B}$ are AZ-ample, we can prove the theorem in a similar way.
In the following, we give an application of the argument of 4.20 for AS-regular algebras. We recall the definitions of AS-regular algebras and AS-Gorenstein algebras below.
@@
The application of 4.20 is the following corollary.